# Surveying the landscape of optically addressable spin qubits for quantum information and sensing technology

### v1

@@ 3 Guidelines for Optimising Performance of Spin-Qubit Materials
From our review, it is clear that spin-based qubit candidates demonstrate potential in several fields of quantum technology. However, unsurprisingly, their spin coherence lifetimes are limited by temperature and spin-bath dependence of the spin properties and the difference in fluorescence between different spin states. Across all material platforms, there is significant magnetic inhomogeneity that emerges from random local spin environments. To understand the extent to which inhomogeneity infects different materials, it could be instructive to consider the ratio of / since 1/ = 1/ + B (Figure 5), where a low ratio indicates that is close to . Using the available data where and were measured under similar conditions, there appears to emerge a distinct advantage for molecules and van der Waals materials despite their lack of isotopic enrichment. This likely stems from the use of molecular crystals and the inherent 2D-order associated van der Waals materials where atomic-scale directional anisotropy ensures that all molecules "feel" the same magnetic environment. By comparison, there are significant difficulties associated with synthesising aligned 3D-spin defects in materials such as SiC and diamond.
@@
Further improvements in spin parameters can be obtained by positional engineering of defect centres. Clustering (or the straggling) of spin-active dopants in a substrate poses spin-spin coupling from the environment (nuclear spins), which decreases . Controlled doping becomes crucial in decreasing spin density around spin-active defect centres. Eliminating unwanted spins like nuclear spins requires isotopic purity of substrate material or the careful doping of spin active centres or defects within the host matrix. More recently, Plasma-Enhanced CVD methods have been used to achieve higher deposition rates while minimising the impact of energetic ions or electrons affecting the colour centres in NV and SnV diamonds[138]. Another novel method to precisely control the position of each defect is laser writing. Aberration-corrected optics allow for the precise positioning of vacancies in diamond systems, with a 45 success probability of a vacancy being located within 200 nm of a desired position[139]. Therefore, chemical systems with charge transfer or hydrogen bonding motifs may yet demonstrate advantages due to their ability to effectively engineer the placement of molecules in 3D. Moreover, as seen with Pc:PTP, and Pr3+:Y2SiO5 vs. the La2(WO5)3 host, materials engineers should avoid defect-site strain by selecting hosts with closely matched physical parameters to the defect. Finally, further improvements can readily be realised through dynamical decoupling methods that are tailored for each application. For example, while engineering clock transitions using magnetic fields or zero-field splitting is appealing for quantum optics where spectral stability is prized, it is not necessarily useful for sensing or information processing. This is because the associated reduced magnetic field sensitivity would reduce the quantum operation fidelity. On the other hand, focused electromagnetic driving of parasitic impurities, such as N-centres in diamond, has yet to be significantly explored in molecular systems. Significant improvements in and would likely emerge from a combination of field-driving and pulsed refocusing methods such as CPMG.
@@ 4 Conclusion and Outlook
Optically addressable electron spin systems show great potential as a diverse form of spin qubit media for quantum sensing, communications, and information processing applications. From this review, we have attempted to benchmark the different families of materials and identify useful investigative and experimental approaches that can be translated across the field. The most significant hurdles faced by chemists and materials engineers remains the strong temperature dependence of the spin-lattice relaxation and thermal polarisation of the spin bath, which renders most d- and f-block metal systems unpracticable above a few kelvin, below which these materials demonstrate the longest coherence times by a significant margin. However, above liquid helium temperatures, it is clear that light-element materials such as colour centres in diamond, SiC and most recently, molecular systems demonstrate more robust quantum spin parameters. Nevertheless, significant advancements in dynamic decoupling techniques within the last 20 years and isotope engineering have enabled the realisation of remarkably competitive quantum spin properties. Within the next decade, it should be expected that further advances and knowledge transfer between investigators of different material platforms with lead to devices capable of significantly impacting society, especially in the field of magnetic field sensors and coherent quantum optics. We hope also that by collating (meta)data on the available spin qubit candidates that the field will eventually benefit from the processing and predictive power of machine learning and artificial intelligence approaches to materials discovery.
@@ 5 Acknowledgements
This work was supported by the UK Engineering and Physical Science Research Council through grants EP/V048430/1 and EP/W027542/1. The authors are also very grateful to Dr. Daan M. Arroo for fruitful discussions and advice.

### v2

@@ 3 Optimising Performance of Spin-Qubit Materials
From our review, it is clear that spin-based qubit candidates demonstrate potential in several fields of quantum technology. However, unsurprisingly, their spin coherence lifetimes are limited by temperature and the spin-bath dependence of the spin properties and optical contrast between different spin states. Across all material platforms, there is significant magnetic inhomogeneity that emerges from random local spin environments. To understand the extent to which inhomogeneity infects different materials, it could be instructive to consider the ratio of / (Figure 10, the so-called "inhomogeneity parameter"). Hence, a low ratio indicates that is close to and the materials are limited by dynamic, not static, sources of decoherence. Using the available data where and were measured under similar conditions, there appears to emerge a distinct advantage for molecules and van der Waals materials at higher temperatures despite their lack of isotopic enrichment. This likely stems from the use of molecular crystals and the inherent 2D-order associated materials such as hBN, which helps to helps to ensure that all molecules "feel" the same magnetic environment. Interestingly, despite the significant difficulties associated with synthesising aligned 3D-spin defects in diamond, SnV-diamond (measured as single spins) exhibits the lowest ratio at low temperatures. Considering its robust quantum spin parameters, SnV-diamond could be a leading candidate for low-temperature applications.
@@
Spin parameters can be improved by positional engineering of defect centres. Clustering (or the straggling) of spin-active dopants in a substrate poses spin-spin coupling from the environment (nuclear spins), which decreases . Controlled doping becomes crucial in decreasing spin density around spin-active defect centres. Eliminating unwanted spins like nuclear spins requires isotopic purity of substrate material or the careful doping of spin-active centres or defects within the host matrix. More recently, Plasma-Enhanced CVD methods have been used to achieve higher deposition rates while minimising the impact of energetic ions or electrons affecting the colour centres in NV and SnV diamonds [170]. Another novel method to precisely control the position of each defect is laser writing. Aberration-corrected optics allow for the precise positioning of vacancies in diamond systems, with a 45 success probability of a vacancy being located within 200 nm of a desired position [171].
Due to their ability to effectively engineer the placement of molecules in 3D chemical systems, materials with charge transfer or hydrogen bonding motifs may yet demonstrate advantages. Moreover, as seen with Pc:PTP, and Pr3+:Y2SiO5 vs. the La2(WO5)3 host, material engineers should avoid defect-site strain by selecting hosts with closely matched physical parameters to the defect. Finally, further improvements can be realised through dynamical decoupling methods that are tailored for each application. For example, while engineering clock transitions using magnetic fields or ZFS is appealing for quantum optics where spectral stability is prized, it is not necessarily useful for sensing or information processing. This is because the associated reduced magnetic field sensitivity would reduce the quantum operation fidelity. On the other hand, focused electromagnetic driving of parasitic impurities, such as N-centres in diamond, has yet to be significantly explored in molecular systems. Significant improvements in and would likely emerge from a combination of field-driving and pulsed refocusing methods such as CPMG.
@@ 4 Conclusion and Outlook
Optically addressable spin systems show great potential as a diverse form of qubit media for quantum sensing, communications, and information processing applications. From this review, we have attempted to benchmark the different families of materials and identify useful investigative and experimental approaches that can be translated across the field. The most significant hurdles faced by chemists and materials engineers remain the strong temperature dependence of the spin-lattice relaxation and thermal polarisation of the spin bath, which renders most d- and f-block metal systems impractical above a few kelvin, below which these materials demonstrate the longest coherence times by a significant margin. However, above liquid helium temperatures, it is clear that light-element materials such as colour centres in diamond, SiC and most recently, molecular systems demonstrate more robust quantum spin parameters. Nevertheless, significant advancements in dynamic decoupling techniques within the last 20 years and isotope engineering have enabled the realisation of remarkably competitive quantum spin properties. Within the next decade, it should be expected that further advances and knowledge transfer between investigators of different material platforms will lead to devices capable of significantly impacting society, especially in the field of magnetic field sensors and coherent quantum optics. We also hope that by collating (meta)data on the available spin qubit candidates that the field will eventually benefit from the processing and predictive power of machine learning and artificial intelligence approaches to materials discovery.
@@ 5 Acknowledgements
This work was supported by the UK Engineering and Physical Science Research Council through grants EP/V048430/1 and EP/W027542/1, and the Department of Materials, Imperial College London. The authors are also very grateful to Dr. Daan M. Arroo for fruitful discussions and advice.
6 Author contributions
CAT and MA curated the data. CAT, MO and MA contributed to the writing of the manuscript.